OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1). parallel: y = -3x + 6 perpendicular: y = x + parallel: y = -3x + 10 perpendicular: y = x parallel: y = x perpendicular: y = -3x parallel: y = x + 1 perpendicular: y = 3x - 1

OpenStudy (anonymous):

???

OpenStudy (anonymous):

if it is perpendicular to a line with slope \(-3\) it will have slope \(\frac{1}{3}\) (the negative reciprocal)

OpenStudy (anonymous):

no need for explanation ... 1 2 3 or 4 ? :( im just trynna finish my test plz :(

OpenStudy (anonymous):

if it goes through the point \((3,1)\) we can write it as \[y-1=\frac{1}{3}(x-3)\] multiply by 3 and get \[3y-3=x-3\] or \[3y=x\]

OpenStudy (anonymous):

i don't see the correct answer from your choices perpendicular line is \(3y=x\) or \(3y-x=0\) so something like that